OpenStudy (anonymous):

Factor: 3y2 - 4y - 15 Which of the following is one of the factors? (3y - 15) (y + 3) (3y + 5) prime

OpenStudy (anonymous):

How do you still not know how to find the answer? You already asked so my questions....you can at least try to figure it out.

OpenStudy (anonymous):

\[3y^2-4y-15=3y^2-9y+5y-15\]\[=(3y^2-9y)+(5y-15)\]\[=3y(y-3)+5(y-3)\]\[=(3y+5)(y-3)\] So 3y+5

OpenStudy (anonymous):

ty